### v1

# TWIN-GPT: Digital Twins for Clinical Trials via Large Language Model

Abstract
Clinical trials are indispensable for medical research and the development of treatments (e.g., drugs). However, a clinical trial often necessitates the involvement of hundreds of participants and can span several years to complete, with a high probability of failure during the process. Recently, there has been a burgeoning interest in virtual clinical trials, which simulate real-world scenarios and hold the potential to significantly enhance patient safety, expedite development, reduce costs, and contribute to the broader scientific knowledge in healthcare. Existing research often focuses on leveraging electronic health records (EHRs) to support clinical trial outcome prediction. Yet, trained with limited clinical trial outcome data, existing approaches frequently struggle to perform accurate predictions. Some research has attempted to generate EHRs to augment model development, but has fallen short in personalizing the generation for individual patient profiles. Recently, the emergence of large language models has illuminated new possibilities, as their embedded comprehensive clinical knowledge has proven beneficial in addressing medical issues. In this paper, we propose a large language model-based digital twin creation approach, called TWIN-GPT. TWIN-GPT can establish cross-dataset associations of medical information given limited data, generating unique personalized digital twins for different patients, thereby preserving individual patient characteristics. Comprehensive experiments show that, using digital twins created by TWIN-GPT can boost the clinical trial outcome prediction, exceeding various previous prediction approaches. Besides, we also demonstrate that TWIN-GPT can generate high-fidelity trial data that closely approximates specific patients, aiding in more accurate result predictions in data-scarce situations. Moreover, our study provides practical evidence for the application of digital twins in healthcare, highlighting its potential significance.
1 Introduction
Clinical trials serve as essential research investigations in the field of prospective medicine, designed to evaluate the safety and efficacy of new treatments (e.g., drugs, or other interventions) rigorously. Despite playing a pivotal role in advancing medical knowledge and enhancing patient care, these meticulously orchestrated studies typically engage tens to thousands of human participants and frequently extend over several years, hindering the progress of research and development. There is an increasing interest in leveraging artificial intelligence models to predict trial outcomes, leveraging electronic health records (EHR) and significantly accelerating the pace of research [21, 56, 23, 33, 19]. Learning from EHR data, models are able to identify the relationship between potential health issues with the personalized characteristics, thus excelling at forecasting disease occurrence and even developing personalized treatment plans [59, 47].
Current approaches leveraging electronic health records (EHR) for clinical trial outcome prediction still face a set of challenges. These methods often fail to account for individual differences between patients and variations among diseases, resulting in predictions that may not align with real-world scenarios [33, 19]. Current EHR-based clinical trial outcome prediction models typically suffers from two key problems: (1) Data Gap: the disconnect between the data and its related realistic background knowledge makes it hard to leverage the supportive knowledge in prediction. (2) Data Inconsistency: discrepancies between data from different sources, presenting different data distributions, different patterns of data missing, and different data recording formats, undermine the reliability of prediction results across datasets.
Recently, digital twin creation technology has exhibited great potential in clinical practice [15, 26], simulating patient physiological state [41, 11], disease progression [76], and treatment effects [1, 13]. These technologies leverages historical and real-time data to provide medical professionals with more accurate and comprehensive personalized health information, introducing opportunities in outcome prediction and analysis. Inspired by this, Das et al[27] designed a simple model for cancer patients’ EHR digital twin creation, which promotes privacy protection and the outcome prediction in clinical trial simulation. However, this method only relies on the existing datasets when creating digital twins, which fails to leverage knowledge beyond the dataset and limits the accuracy of clinical trial outcome prediction. Traditional digital twin models still face challenges in achieving accurate clinical trial outcome predictions, failing to effectively address data gaps and data inconsistencies. Data gaps are particularly prominent in trial predictions due to variations in ICD codes for different medications and diseases, necessitating the acquisition of real-world background knowledge by the digital twin for clinical trials. Moreover, the data inconsistency problem is greatly exacerbated in EHR data for clinical trial scenarios, posing a challenge to digital twin creation, since EHRs for clinical trials often originate from different sites and thus present greater inconsistencies.
To address these problems, in this paper, we propose a novel approach based on LLMs that creates digital twins to enhance the effectiveness and accuracy of clinical trial outcome prediction. Often referred to as ‘world models’, LLMs possess not only extensive language understanding capabilities but also a comprehensive understanding of world knowledge, including medical knowledge. This makes LLMs capable of overcoming the aforementioned data gap and inconsistency problems. Our proposed innovative model TWIN-GPT is fine-tuned on a pre-trained LLM (ChatGPT [53]) on clinical trial datasets, so as to generate personalized digital twins for different patients. According to the virtual personalized patient data (i.e., digital twins) generated by TWIN-GPT based on historical and real-time examination results, we find that patients’ physiological and pathological states over the clinical trial duration can be accurately forecasted, thereby achieving more precise trial outcome predictions. Moreover, experiments on real-world datasets demonstrate that our TWIN-GPT can account for individual patient variations and disease complexities, producing data that closely aligns with diverse real-world scenarios. It resolves challenges of data gap and inconsistency in EHR while ensuring the protection of patient privacy. Our main contributions are summarized as follows:
  • $\bullet$
    Innovative Integration of LLM for Digital Twin Creation: To the best of our knowledge, we are the first to integrate LLM into digital twin creation and perform knowledge association across datasets, which effectively imputes missing EHR data and provides a more personalized patient modeling approaches, thus addressing the limitations of traditional models.
  • $\bullet$
    Enhanced Personalization and Accuracy: TWIN-GPT harnesses the vast medical knowledge embedded within ChatGPT to generate personalized digital twin models for individual patients. Experimental findings showcase that this approach achieves significantly enhanced personalization by accounting for each patient’s unique characteristics and disease complexities, thereby improving the accuracy of predicted clinical trial outcomes.
  • $\bullet$
    Privacy Protection and Versatility in Application: Our TWIN-GPT approach also protects the patient privacy by generating virtual patient data and simulating personalized physiological measures over time, minimizing the use of sensitive patient information. The versatility allows for application across various clinical trial scenarios, thereby accelerating the pace of clinical trials and enhancing medical research and patient care [27].
2 RELATED WORK
2.1 Large Language Models (LLMs) for Medicine
Recently, LLMs have exhibited excellent performance on various tasks, especially in the field of medicine [62, 25, 45]. The current research on LLMs in the medical field has provided novel applications in bedside diagnosis, automated drafting of clinical documents, and improving medical workflow efficiency [12, 54, 51]. LLMs have demonstrated their utility in clinical applications by passing medical licensing exams and offering superior empathy and quality of medical advice compared to human clinicians [69, 42, 52, 61, 71, 75, 73]. In clinical applications, LLMs were used to facilitate the analysis of large amounts of textual data, providing insights that can drive medical discoveries and increase the efficiency of research workflows [66, 67, 60, 4]. They also have significant advantages in daily clinical tasks like clinical documentation drafting, and can significantly reduce the administrative burden on medical staff [78, 36, 10, 30]. However, there is no research using LLMs in clinical trial scenarios, which is a crucial area demanding the extensive medical knowledge encoded within these large language models.
2.2 Patient Outcome Prediction
Previous research in patient outcome prediction primarily relied on traditional machine learning methods such as decision trees, support vector machines, and logistic regression [20, 40, 14]. These approaches often utilized manually selected features from high-dimensional clinical data. However, manual feature selection frequently fails to incorporate informative features, thereby failing to fully capture individualized patient information and complex data relationships [22]. Recently, deep learning methods have gained prominence with the advancement of deep learning technology. Researchers have begun to explore the use of neural networks for patient prediction [21, 56, 23, 44], which handle well with large-scale data and automatically select informative features, leading to significant achievements. Fu et al. [33, 32] proposed a hierarchical interaction network (HINT) to predict the clinical trial outcome based on drug molecules, disease codes, eligibility criteria, etc. Lu et al. [48] enhance HINT from the perspective of uncertainty quantification and interpretability.
Among these, analysis of EHR data plays a crucial role in patient outcome prediction. Many studies have investigated how to effectively utilize EHR data for patient prediction [20, 39, 55], applying long-term historical EHR learning, medical text mining, time-series data analysis [43, 2, 79]. Utilizing EHR data can provide comprehensive patient information, but it also suffer from data gap problem and data inconsistency problem. Recently, digital twin creation has emerged as a noteworthy field in EHR learning, offering new opportunity to patient outcome prediction [63]. Digital twin models can simulate a patient’s personalized physiological states, disease progression, and treatment effects based on historical EHR and real-time data [27].
2.3 EHR Data Generation
Data generation methods can create synthetic data that is effective in addressing data scarcity issues. EHR data generation methods often synthesize patient records conditioned on a patient’s historical data and distribution information, using techniques such as generative adversarial networks (GANs) [5, 22, 17] or variational auto-encoders (VAEs) [1, 7, 18, 34, 31]. The advantage of these methods is their ability to protect patient privacy, but ensuring the generated data maintains consistency with real data [57] remains a challenge.
To protect patient privacy, many approaches employ data masking and anonymization techniques in learning EHR data [35, 77, 65, 68]. These methods involve de-identifying or anonymizing sensitive information to create anonymous or partially anonymous data. While these methods provide privacy protection, they may potentially lead to a decrease in data quality [28, 8]. Model-driven generation methods use known patient models or epidemiological knowledge to generate synthetic EHR data. These models can take into account a patient’s physiological states, disease progression, and treatment effects, resulting in more accurate data generation [80, 9, 24]. Typically, these methods require domain experts’ knowledge to guide the generation process. Some data augmentation techniques can also increase data diversity and quantity using existing EHR data. This includes extracting information from medical texts using natural language processing techniques or creating new data points by combining different types of medical events [16, 72, 29, 74]. Data augmentation can enhance the training effectiveness of models but is still subject to limitations imposed by the original data.
Digital twin is a specialized method for generating personalized EHR data, ensuring similarity and correspondence with the original data, as well as guaranteeing patient privacy protection [49]. Das et al[27] used a VAE to model digital twins in clinical trials to achieve prediction of patient outcomes in low-data scenarios. Liu et al[46] proposed a cloud healthcare system framework based on digital twin healthcare (CloudDTH), which is used to monitor, diagnose, and predict personal health to achieve the goal of personal health management. Zhong et al[81] used ICU EHR data to build the ICU digital twin model to study critical care services in the ICU. However, these approaches only perform digital twinning within their datasets and do not solve the data gap and data inconsistency problems.
3 METHODOLOGY
In this section, we introduce TWIN-GPT, a novel approach that creates personalized digital twins for effective virtual clinical trial simulation, leveraging the power of large language models (LLMs) to address the challenges of clinical trial outcome predictions in limited data scenarios. Our method is predicated on the understanding that clinical trials are critical for medical research and drug development but are hampered by the extensive time and participant involvement required. TWIN-GPT aims to revolutionize this by generating high-fidelity, personalized digital twins of patients, thus facilitating virtual clinical trials that can significantly enhance patient safety, expedite development, and reduce costs.
In Section 3.1, we define the problem of creating a digital twin model for clinical trials. Section 3.2 details the intricacies of data input encoding and the data preparation necessary for effective model training. In Section 3.3, we introduce training strategies for synthesizing digital twins from patient data. Then, Section 3.4 elaborates on the training process for the digital twin model, enhancing its adaptability to individual patient data. Finally, Section 3.5 discusses the generation of digital twins, illustrating how the model manipulates synthesized data to simulate realistic EHR data.
3.1 Problem Formulation
We assume that there are a total of $N$ participants in a clinical trial. Each participant $n$’s clinical encounters are encapsulated as
$X_{n;1:T_{n}}=\left\{x_{n,1},x_{n,2},\dots,x_{n,T_{n}}\right\},$ (1)
where each $x_{n,t}$is a collection of event types occurring at the $t$-th visit, further delineated as
$x_{n,t}=\left\{x_{n,t}^{1},x_{n,t}^{2},\dots,x_{n,t}^{U}\right\},$ (2)
$x_{n,t}$ capturing the diverse clinical events ranging from treatments to adverse reactions. where $U$ is the total number of all types of events. An event set $x_{n,t}^{u}$ contains all u type events that occurred at the $t$-th visit for the patient $n$. $x_{n,t}^{u}=\left\{c_{1},c_{2},\dots,c_{n}\right\}$ represents all events of type u, and the value of $c_{l}$ is 1 or 0, indicating whether the events happened or not. Utilizing TWIN-GPT, we synthesize digital twins $\hat{X}_{n;1:T_{n}}$ by generating personalized patient trajectories based on their real-world counterparts and nearest neighbor patients, encapsulating a wide spectrum of medical conditions and treatments. Formally, this is represented as:
$\hat{X}_{n;1:T_{n}}=\text{TWIN-GPT}\left(X_{n;1:T_{n}},\{X_{n^{\prime};1:T_{n^{\prime}}}\}_{n^{\prime}=1}^{N^{\prime}},\Theta\right),$ (3)
where $\Theta$ denotes the learned model parameters, and $N^{\prime}$ is the count of nearest neighbor patients considered for synthesizing the digital twin. Then we identify the most similar feature vector from the dataset, which will serve as the ground truth:
$GT_{n,t}={sim}\left(\bar{F}_{n,t},\{F_{n^{\prime};t^{\prime}}\}\right),$ (4)
where $GT_{n,t}$ represents the ground truth for patient $n$ at visit $t$, derived from the pool of averaged feature vectors $\{F_{n^{\prime};t^{\prime}}\}$ across all patients and visits.
In our analysis, we particularly focus on three critical aspects: treatment approaches, medication usage, and adverse events. Treatment is the assigned treatment at the $n$-th timestep for the participant $n$, denoted $x_{n,t}^{\text{treat}}$. medication denoted as $x_{n,t}^{\text{med}}$, is defined as the treatment drugs given to patients in response to their current health conditions. Adverse events represented by $x_{n,t}^{\text{ae}}$, refer to all unexpected health incidents that occur during patient visits. To let the model predict the adverse events in the next visit as $X_{n,t+1}^{\text{ae}}$, we choose to input all three vital type events represented by $X_{n,t}^{\text{treat}}$, $X_{n,t}^{\text{med}}$ and $X_{n,t}^{\text{ae}}$ because the causality between three types of events play an important role in predicting. For example, the doctor might provide medication Docetaxel to the patient at timestep $t+1$ because of the adverse events observed in the last visit. Meanwhile, the adverse event pain might not be observed at timestep $t+1$ due to the medication Docetaxel given at timestep $t$. It is easy for large language models, such as ChatGPT, to fully capture the causality between all kinds of events because they have robust in-context learning ability [50].
3.2 Input Encoding
In the process of clinical trials, we define the original patient record $X_{n;1:T_{n}}$ by collecting the sequence of patient visit records. Our goal is to generate a digital twin $\hat{x}_{n;1:T_{n}}$ that can retain the features of the target patient. For this purpose, we propose a generator, namely TWIN-GPT, that simulates personalized patient trajectories, which are determined by the real record $X_{n;1:T_{n}}$, $\{X_{k;1:T_{n}}\}_{k}$ is the nearest $K$ neighbor patients, and $\theta$ is the learned model parameters that are constantly updated during training, as shown below:
$\hat{X}_{n;1:T_{n}}=\textit{TWIN-GPT}(X_{n;1:T_{n}},\{X_{k;1:T_{k}}\}_{k};\theta).$ (5)
Each visit $x_{n,t}$ within the primary input $X_{n;1:T_{n}}$ is structured from $U$ diverse concurrent event types, illustrated as $x_{n,t}=\{x^{1}_{n,t},\ldots,x^{U}_{n,t}\}$, forming a chronological sequence of visits $x_{n,t}$.The event set $x^{u}_{n,t}$ is converted into a multi-hot vector, where each element represents the occurrence of a specific event at this time step, i.e., $x^{u}_{n,t}=\{c_{1},\ldots,c_{l}\}$, where $c_{i}\in\{0,1\}(i=1,2,3,...,l)$.
The initial input $X_{n;1:T_{n}}$ is constructed from a series of patient visits $x_{n,t}$, each visits $x_{n,t}$ encompasses $U$ distinct categories of events simultaneously, such as $x_{n,t}=\{x^{1}_{n,t},\ldots,x^{U}_{n,t}\}$. The event set $x^{u}_{n,t}=\{c_{1},\ldots,c_{l}\}$ is converted into a multi-hot vector, where each element represents the occurrence of a specific event at this time step, where $c_{i}\in\{0,1\}(i=1,2,3,...,l)$.
3.3 Prompt Tasks
The prompt task is a piece of text input provided to the model to generate a response based on its training. We proposed a unique approach to train Electronic Health Record (EHR) model. This approach is based on individual patient sessions and generates a digital twin of the patient through a series of steps. The process includes the following stages:
(1) Initially, a digital twin is established based on the patient’s EHR data. At this stage, the patient’s EHR data and five EHR datasets similar to the patient’s condition are provided. These data are input into the process of generating the digital twin. (2) Next, the correct answer is provided, with the expectation that the training process will be optimized by summarizing and generalizing these answers. This process is iterated according to the number of patient visits. For each visit, the patient’s profile and the EHR data from the first visit are provided, followed by five EHR profiles similar to the patient’s symptoms. Based on this, an adverse event that the patient may experience under the current treatment is predicted. If the predicted outcome deviates significantly from the correct answer, multiple-choice answers will be provided [37]. (3) Subsequently, the correct adverse event under the current treatment is provided. Based on these adverse events and the treatment, the medication that the patient should use next time is deduced and this information is provided. (4) Finally, a digital twin of the patient’s medication and adverse events is generated. This digital twin must take into account the specific treatment name.
This process is repeated each time the patient’s EHR data is updated. The updated EHR data is provided, along with five other EHR profiles similar to the patient’s symptoms, and then adverse events are predicted and medication is recommended again. The goal of this approach is to track and update the patient’s EHR digital twin in real-time, in order to provide more accurate treatment recommendations.
Figure 1: The workflow of TWIN-GPT. (Bottom) TWIN-GPT takes real follow-up visits $X_{n,1:T_{n}}$ of a patient and generates twin visits, $\hat{X}_{n,1:T_{n}}$. (Top) TWIN-GPT uses digital twin $\hat{x}_{n,t}$ to predict the events that occurred in the next timestamp $\hat{x}^{\text{event}}_{n,t+1}$ using a treatment prediction related prompt. We use K nearest neighboring visits for fine-tuning but only use origin visits in prediction. “Treat”: treatment; “Med”: medication; “AE”: adverse event.
3.4 Digital Twin Modeling
To enable TWIN-GPT to learn personalized clinical trial data for patients and to realize data association between different patients, we adopted the method of prompt fine-tuning. In this way, TWIN-GPT is able to accurately digitally twin the characteristics of the patient, and finally realize the prediction of different clinical trials.
Firstly, we fine-tune TWIN-GPT by letting it make predictions based on the patient’s historical access logic so that TWIN-GPT can learn the patient’s access pattern. That is, by predicting the drug treatment and treatment plan at the current time point $t$, the Adverse Event (AE) at the next time point $t+1$ can be predicted so that TWIN-GPT can improve the prediction accuracy of future adverse events by learning the access pattern and drug treatment information in historical data. In particular, in the prediction of AE, we selected $n$ nearest neighbors by the $k$-NN method and processed them accordingly.
At the same time, to realize the association of TWIN-GPT with patient data at different access time points, we fine-tune it to learn the relationship between the data at different time points, better capture the dynamic changes of patient data, and improve TWIN-GPT’s comprehensive and comprehensive understanding of individual patient history. Specifically, we assume the twin process is reversible. That is, there is a mapping relationship that allows you to map the source data space to the twin data space and the twin data space to the source data space.
For each visit $x_{n,t}$, we search all the other patients we retrieve $K$ visits $\left\{x_{k,t^{\prime}}\right\}^{K}_{k}$ from all the other patients with the highest similarity to $x_{n,t}$. Note that $t^{\prime}$ can be either equal or unequal to $t$. Then, we calculated a $x^{\prime}_{n,t}$ based on their similarity, where each number represents the probability of the event occurring. We set $x^{\prime}_{n,t}$ as the input and the $x_{n,t}$ as the target for TWIN-GPT in the twin fine-tuning process. We also fine-tuned the next visit prediction using each patient’s next visit event list, i.e., for the target patient $n$ at visit $t$, using his medication events $x^{\text{med}}_{n,t}$ and treatment events $x^{tr}_{n,t}$ as input and his next adverse events $x^{\text{ae}}_{n,t+1}$ as label for TWIN-GPT in the prediction fine-tuning process. The two fine-tuning processes are carried out alternately so that TWIN-GPT can learn the relationship of all kinds of events.
3.5 Digital Twins Generation
3.5.1 Personalized Generation
Personalized Generation refers to the process of using EHR data to guide TWIN-GPT in replicating patients, ultimately generating digital twins. Throughout each step of this process, the model incorporates both its data and data from the k-nearest neighbors, enabling TWIN-GPT to engage in in-context learning.
The workflow in Fig. 1 can be summarized in the following key steps: Firstly, we acquire the patient’s self-data (visit) from their EHRs. Next, we perform calculations to determine the $k$-nearest neighbors for this visit, which includes important components such as MED and AE. Subsequently, we apply a masking operation to the self-visit data and design specific prompts for TWIN-GPT. These prompts have two primary tasks: reconstructing the masked data portion and predicting the current treatment strategy.
TWIN-GPT takes over the process, performing both data reconstruction and treatment prediction tasks, ultimately yielding a virtual entity known as “visit:n”. Following this, we provide TWIN-GPT with ground truth information regarding the treatment, enabling it to perform self-corrections and make predictions for the next visit ($visit_{n}+1$).
3.5.2 Counterfactual Generation
The purpose of this task is to simulate the trajectories of patients under alternative treatment schedules, including switching a patient from the treatment arm (T) to the control arm (C). This simulation not only enhances patient records but also enables the estimation of personalized treatment effects. Moreover, it facilitates the balancing of trial data for predictive modeling, while substantially reducing the necessary sample size for recruiting participants in the control arm. To generate the counterfactual digital twin $\hat{X}_{n,1:T_{n}}$, corresponding to the real patient record $X_{n,1:T_{n}}$, we need to identify the most closely matching patient record within $\tilde{X}_{k,1:T_{k}}$, where $n$ is among $T$ and $k$belongs to $C$. This step merges the unique personal traits of patient $n$ with the temporal patterns observed in patient $k$, facilitating the creation of the synthetic pathway. We evaluate the applicability of TWIN-GPT trajectory generation by comparing the similarity between the synthesized trajectory and the most similar patient record.
4 EVALUATION METRICS
We evaluate the generated data of digital twins based on EHR data from two aspects, aiming at clinical trial outcome prediction and generative privacy.
4.1 Clinical Trial Outcome Prediction Evaluation of Digital Twins
To assess the practicality of models, we observe whether the additional synthetic data generated by models can successfully predict outcomes under different treatment modalities. Specifically, our experiments involve the following prediction tasks:
4.1.1 Dimension-wise probability.
To evaluate the similarity between models’ twin data and the real data, we perform a dimensional-wise probability calculation. The calculation of dimension probabilities helps us to understand the probability of occurrence of each feature in the dataset, and it can provide quantitative metrics about the consistency and similarity between synthetic and real data. The Dimension-wise probability (DP) is calculated by
$DP=\frac{V_{f}}{V_{l}},$ (6)
where $V_{f}$ is the number of visits in the dataset containing a specific feature. $V_{l}$ is the total number of accesses in the dataset. When the model successfully produces synthetic data that closely mirrors the original, the distributional properties (DPs) of the synthetic data will approximate those of the real data. We calculate the Person correlation coefficient, $r\in[0,1]$. If $r=1$, it means that the model has high fidelity.
4.1.2 Counterfactual digital twin evaluation.
In the evaluation of counterfactual digital twins, for a specific patient, we identify the closest match from different treatment groups to act as a substitute for the inaccessible counterfactual outcomes, in line with established practices in causal inference research [58]. We assess the similarity between the digital twin and its corresponding record by calculating personalized Pearson correlation coefficients and quantify the fidelity by comparing data points between the synthetic and surrogate records.
4.1.3 Severe outcome prediction
To validate whether TWIN-GPT can fully comprehend and replicate the underlying relationship between real-world data records and severe outcomes, we conducted severe outcome prediction validation. Firstly, we defined severe outcomes to include deaths and other critical clinical events. Then, we employed Long Short-Term Memory (LSTM) to forecast these severe outcomes. LSTM received real clinical data and synthetic data generated by different models for model training and prediction. Ultimately, we utilized the Area Under the Receiver Operating Characteristic Curve (AUROC) to assess the accuracy and performance of the predictions.
4.1.4 Adverse event prediction
Adverse event prediction is to see how much causal relation the synthetic data generated by TWIN-GPT have maintained compared to the raw clinical data. We train another MLP to evaluate the prediction performance, by inputting synthetic data and real data separately.
4.2 Digital Twins Clinical Privacy Evaluation
In the generation and utilization of synthetic data, privacy protection is of paramount importance. This section will delve into three primary methods of privacy risk evaluation: Presence Disclosure, Attribute Disclosure, and Nearest Neighbor Adversarial Accuracy Risk (NNAA).
4.2.1 Presence Disclosure
To evaluate the risk of privacy leakage, where an attacker analyzes synthetic data to infer whether the records of a particular individual participated in the training of the model, we employ Sensitivity as a crucial metric. Sensitivity serves as an important indicator for assessing the privacy risk associated with synthetic data. By comprehending sensitivity, we can gain a clearer understanding of the level of privacy protection offered by synthetic data and implement appropriate measures to mitigate the risk of existential leakage. The sensitivity is calculated using the following formula:
$S=\frac{R_{k}}{R_{l}},$ (7)
where $R_{k}$ is the number of known records found in the synthetic data and $R_{l}$ is the total number of known records. For instance, if we have 100 known records and detect 20 of them in the synthetic data, the sensitivity would be $20/100=0.2$. This implies a 20% probability for an attacker to confirm the presence of known samples in the training data. By calculating the sensitivity, we can determine the likelihood of an attacker identifying a known sample within the training data. If the synthetic data contains a limited number of known records, resulting in low sensitivity, it indicates a relatively secure nature of the synthetic data with a reduced risk of existential leakage. Conversely, if the synthetic data contains a substantial number of known records, leading to a high sensitivity, the risk of existential leakage becomes more pronounced.
4.2.2 Attribute Disclosure
Privacy leakage is when an attacker is able to obtain or infer sensitive information or unknown attributes of an individual, thereby violating an individual’s right to privacy. Attribute leakage occurs when the attacker is able to infer other attributes of the patient based on a known subset of the data. To quantify this risk, we adopt average sensitivity as a metric. Average sensitivity measures the attacker’s ability to infer unknown attributes. It is calculated using the following formula:
$MS=\frac{1}{N}\sum_{v=1}^{N}\frac{F_{d}}{F_{l}},$ (8)
where $F_{d}$ represents the number of unknown features discovered, and $F_{l}$ represents the total number of unknown features. For example, if an attacker knows 10 attributes of a patient and can infer 2 unknown attributes from the synthetic data, the average sensitivity would be $2/10=0.2$. This implies that the attacker has a 20% probability of confirming the presence of these unknown attributes in the synthetic data. By computing the average sensitivity, we gain insight into the attacker’s average capability to infer the unknown attributes of the patient. A higher average sensitivity indicates a greater risk of privacy leakage, as it suggests that the attacker has a higher likelihood of inferring unknown attributes. Conversely, a lower average sensitivity indicates a lower risk of privacy leakage, as it implies that the attacker’s inference capability is relatively weak.
4.2.3 Nearest Neighbor Adversarial Accuracy Risk (NNAA)
NNAA is a measure of the degree to which a model overfits the original data, directly relating to privacy leakage risk. The NNAA risk score is calculated as follows:
NNAA risk $\text{NNAA risk score}=AA_{ES}-AA_{TS},$ (9)
where $AA_{ES}$ and $AA_{TS}$ represent the aggregated distances between synthetic data and evaluation data, and between synthetic data and real data, respectively. They are calculated as follows:
[…]
Here, $d_{ES}(i)$, $d_{TS}$, $d_{SE}$ and $d_{EE}$, $d_{TT}$, $d_{SS}$ represent the nearest neighbor distances among synthetic data, evaluation data, and real data respectively. For example, if the average distance between synthetic data and evaluation data is greater than that between synthetic data and real data, the NNAA risk score will be positive, indicating potential overfitting of the model to the training data, hence increasing the risk of privacy leakage.
Through these three methods, we can evaluate the privacy risks of synthetic data more comprehensively, thereby better protecting individual privacy.
5 EXPERIMENTS
5.1 Experimental Setup
5.1.1 Data source
Original clinical trial dataset: We trained TWIN-GPT using a Phase III breast cancer trial dataset (NCT00174655), analyzing Disease-Free Survival (DFS) across different treatment methods. The study randomly divided 2,887 patients into groups to compare the efficacy of docetaxel, either alone or with doxorubicin followed by CMF, against doxorubicin alone or combined with cyclophosphamide, then CMF, in patients with positive axillary lymph nodes. This dataset, publicly available in Project Data, facilitated our performance evaluation. Sphere [^1] https://www.projectdatasphere.org. Trial Outcome Prediction (TOP) dataset: We also utilized the multi-modal TOP dataset [32] for predicting clinical trial outcomes. This dataset integrates clinical trial-related information from multiple data sources, with each trial record including drug molecule information, disease information, eligibility criteria, and trial result information. The dataset is divided into three phases: 1,160 Phase I trials, 4,449 Phase II trials, and 3,436 Phase III trials. The drug molecule information includes the names of candidate drugs and their functional groups. The disease information includes the ICD-10 codes, disease descriptions, and disease hierarchy represented by CCS codes. The eligibility criteria for the trials are described using unstructured natural language, including inclusion and exclusion criteria. The trial result information includes binary indicators for trial success (1) or failure (0), trial phase, start and end dates, trial sponsor, and trial scale (i.e., the number of participants).
5.1.2 Data preprocessing
We processed this dataset by extracting medications, treatments, and adverse events from the original clinical trial data. Then we selected the top 100 most frequently used medications and the top 50 most frequent adverse events, which contain most of the information in this dataset. We also merged all rare events (frequency < 50) into one additional adverse event, so that the model could fully utilize all samples. Regarding the TOP dataset [32], we utilized three distinct phases of this dataset to construct a twin network model. The study fully capitalized on the complete information available in the TOP dataset, wherein the criteria for trial qualification were delineated as background knowledge for establishing twin states, whereas the rest was considered clinical medical data for developing twin characteristics.
5.1.3 Baseline Models
We compare TWIN-GPT with the following baseline methods:
  • EVA [7] utilizes variational autoencoders to create synthetic versions of electronic health records, serving as a generative model for this purpose.
  • SynTEG [80] is a representative generative model that employs GANs to produce synthetic versions of EHRs. It uses transformer [64] structure in the dependency extraction part and Wasserstein GAN with gradient penalty (WGAN-GP) [3, 38] in the conditional generation part.
  • PromptEHR [70] leverages real EHRs to train a prompt learning-based generative language model for synthetic EHR generation.
  • TWIN-VAE [27] can detailed clinical data through the development of individualized clinical trial digital twins, utilizing variational autoencoders (VAE) for this process.
  • $k$-NN-based method [6] is a simple model that modifies the real patient data by picking random parts from its nearest neighbors.
5.2 Generation Quality
5.2.1 Personalized Generation.
To make sure that TWIN-GPT has fully learned every dimension’s distribution precisely and accurately, we calculate all DPs of every adverse event separately from both real and generated synthetic data. Fig 3 shows the performance of TWIN-GPT for Breast Cancer data. We compare the performance of dimensions’ distribution of TWIN-GPT and other baselines. The rs of adverse events for EVA, SynTEG, PromptEHR, $k$-NNand TWIN-VAE are -0.06, -0.07, 0.14, 1.0 and 0.99, respectively [27]. We can see that TWIN-GPT has higher fidelity than SynTEG, PromptEHR, and TWIN-VAE. Meanwhile, $k$-NN achieves performance comparable to TWIN-GPT, which can be easily explained that $k$-NN essentially copies and merges fragments from real data to generate synthetic data. However, $k$-NN cannot be used to generate EHR data directly because all the generated data can be linked to their original patient’s record, which will lead to the privacy risk of leaking information. Regarding the TOP dataset, we conducted overall DP calculations for the three medical phases in Fig 3 and independent DPs calculations for the three clinical trial phases as shown in Fig 4. In both scenarios, TWIN-GPT demonstrated high performance with $\geq$0.96. Since the TOP dataset is a multimodal dataset, the comparison algorithms used in the previous dataset do not apply to this dataset. Here, we used $k$-NN to calculate for the TOP dataset, where its $r$ is 1, slightly higher than our TWIN-GPT. As mentioned before, $k$-NN is not suitable due to the potential for privacy risks associated with linking generated data back to the original patient records.
Figure 2: On the Original clinical trial dataset, we analyzed the dimension-wise Pearson correlation coefficient (r) of adverse events to evaluate the performance of TWIN-GPT. The x-axis displays the probability across dimensions for real data, while the y-axis signifies the probability associated with synthetic data.
Figure 3: On the TOP dataset, we analyzed the dimension-wise Pearson correlation coefficient (r) of adverse events to evaluate the performance of TWIN-GPT. The x-axis displays the probability across dimensions for real data, while the y-axis signifies the probability associated with synthetic data.
(a) Phase I
(b) Phase II
(c) Phase III
Figure 4: On the TOP dataset, we analyzed the dimension-wise Pearson correlation coefficient (r) of adverse events in the three phases to evaluate the performance of TWIN-GPT.The x-axis displays the probability across dimensions for real data, while the y-axis signifies the probability associated with synthetic data.
We also find out per-patient DPs to determine how well each synthetic record matches the associated synthetic record. We compute the Pearson Correlation Coefficient, r. The comparison of distributional properties (DPs) between synthetic and real data across all patients is analyzed through Pearson Correlation Coefficients, shown in a histogram within Fig. 6. This comparison indicates the synthetic data’s high fidelity on a feature-wise level, with a significant portion of correlation coefficients (r values) surpassing 0.8.
5.2.2 Counterfactual Generation.
To assess the quality of the counterfactual generation results, we train an LSTM model on real data to predict severe outcomes. Subsequently, we utilize this trained model to predict the counterfactual digital twins generated by TWIN-GPT, representing the simulated patients assigned to the alternative treatment arm. We then compare these predictions with the surrogate ground truth outcomes. Notably, the obtained AUROC score of 0.821, which is quite close to 0.838, AUROC score of the predictions with real data, demonstrates the high fidelity of the generated counterfactual digital twins.
Just like the way we determine the feature-wise fidelity in the Personalized Generation part, we assess the distributional properties (DPs) of actual data from the closest neighbor against the DPs of the corresponding synthetic data to determine the Pearson correlation coefficient ($r$). This comparison yields coefficients that are represented in Fig 6. The results indicate that the digital twins of most patients exhibit a high degree of similarity to their real-life counterparts. These digital twins exhibit high $r$s that are larger than 0.8.
Figure 5: Patient-wise Pearson correlation coefficient (r) for TWIN-GPT. r is charting the distributional properties (DPs) of each patient’s closest match on the x-axis against the DPs of their respective synthetic digital twin on the y-axis. Most of the participants have high fidelity (r larger than 0.8).
Figure 6: Counterfactual Generation (r) for TWIN-GPT. r is charting the distributional properties (DPs) of each patient’s closest match on the x-axis against the DPs of their respective synthetic digital twin on the y-axis. Most of the participants have high fidelity (r larger than 0.8).
5.3 Digital Twins Clinical Trial Prediction
5.3.1 Severe outcome prediction
We train an LSTM to predict the severe outcome. This model takes in the real clinical data and synthetic data generated by TWIN-GPT. We show this severe outcome prediction result in Fig 8. The LSTM achieves nearly the same AUROC scores when taking in the synthetic data generated by TWIN-GPT as taking in the real data, indicating that TWIN-GPT can fully understand and reproduce the latent relationship between records and severe outcome in real data.
Figure 7: Severe outcome prediction results measured by AUROC scores. The AUROC scores of prediction made by LSTM when taking in real and synthetic data are quite close.
Figure 8: Adverse event prediction AUROC scores. The red line means the AUROC score of MLP to predict the adverse event by taking in real data. We can see that the synthetic data generated by TWIN-GPT and $k$-NN-based model have similar performance when the MLP predictor is trained on synthetic data and real data.
5.3.2 Adverse event prediction
This task is to see how much causal relation the synthetic data have maintained compared to the raw clinical data. We train an MLP to predict the adverse event at the next time step, by taking in real and synthetic data generated by different models separately. Results are shown in Fig 8, where both TWIN-GPT and $k$-NN-based model have similar performance with the real data, indicating that our method captures the temporal causal relations accurately.
5.4 Digital Twins Clinical Privacy Protection
Privacy protection is of paramount importance in the generation and utilization of synthetic data. In this part, we evaluate three primary indexes: Presence Disclosure, Attribute Disclosure, and Nearest Neighbor Adversarial Risk (NNAA).
(a) Presence disclosure
(b) Attribute disclosure
Figure 9: On the Original clinical trial dataset, Presence disclosure and Attribute disclosure sensitivity scores with a different number of samples known by the attacker. (Lower sensitivity is better).
(a) Presence disclosure: comprehensive perspective
(b) Presence disclosure: phase-by-phase analysis
Figure 10: On the TOP dataset, presence disclosure sensitivity scores with a different number of samples known by the attacker. Lower sensitivity is better.
(a) Attribute disclosure: comprehensive perspective
(b) Attribute disclosure: phase-by-phase analysis
Figure 11: On the TOP dataset, Attribute disclosure sensitivity scores with a different number of samples known by the attacker. Lower sensitivity is better.
5.4.1 Presence Disclosure
If an attacker finds that the synthetic data was trained by TWIN-GPT from the patient n’s record $X_{n;1:T_{n}}$, we call it presence disclosure. We assume that there are total $m$of training data that had been known by the attackers, and set $m$ as 1%, 5%, 10%, and 20%. After calculating the sensitivity through Eq.7, we display the result in Fig 9(a). As the number of known samples rises, the sensitivity of TWIN-GPT has remained stable at around 20% (the attackers can recognize 20% of the real patients’ records from synthetic data they know). Correspondingly, the $k$-NN-based method reaches the maximum sensitivity of around 50%. Additionally, we observe that in terms of presence disclosure, TWIN-GPT’s sensitivity is slightly higher than that of TWIN-VAE but significantly lower than the $k$-NN algorithm. This could be because TWIN-GPT, as a large language model-based method, might possess a greater capability to capture and replicate the subtle differences of the training data. Such high fidelity in data replication may inadvertently lead to synthetic data retaining too many identifiable features of the original training data, thus increasing the risk of presence disclosure. However, on the TOP dataset, both the overall medical trial sensitivity for presence disclosure and the phase-by-phase medical trial sensitivity for presence disclosure are better than $k$-NN in Fig 10.
5.4.2 Attribute Disclosure
Here, we assume that the attackers know parts of the training set and have access to $x$of features of the records. We set $x$ as 1, 5, 10, and 20 in this experiment. When an attacker can infer additional attributes of an individual based on the features of a subset of the data he knows, the attribute disclosure occurs. We use Eq.8 to calculate mean sensitivity. The results are shown in Fig 9(b). We can see that the mean sensitivity of TWIN-GPT is generally less than the $k$-NN-based method and Twin method, meaning better performance. We can also observe that the maximum score of $k$-NN-based method and TWIN-GPT are respectively around 0.3 and 0.25. On the TOP dataset, as shown in Fig 11, the mean sensitivity of TWIN-GPT is significantly lower than that of the $k$-NN algorithm. Across different numbers of features, TWIN-GPT’s mean sensitivity is 0.2 percentage points lower than that of $k$-NN.
5.4.3 Nearest neighbor adversarial accuracy risk
NNAA is a measure of the degree to which a model overfits the original data, directly relating to privacy leakage risk. Here, we select 71 participants with 500 visit records as evaluation set $S_{E}$. Correspondingly, we also choose 500 clinical records from real data as training set $S_{T}$ and from synthetic data to form $S_{S}$, like the method required. TWIN-GPT achieves a score of 0.271. Generally, If the NNAA score is close to 0.5, we say that models overfit the original data. Therefore it suggests that TWIN-GPT is not excessively memorizing the original data and is instead learning more generalizable patterns.
5.5 Model Explainability
Unlike traditional models, our model functions as a “world model” endowed with a rich set of capabilities, enabling it to understand and interact with complex data landscapes. Therefore, our model possesses strong explainability. To this end, we conducted the explainability test of the twin generation results. As shown in Fig. 12, TWIN-GPT can provide explanations for its results when generating twins from input data.
Figure 12: Explanation of TWIN-GPT.
5.6 Ablation Study
We compare the performance of TWIN-GPT before prompt fine-tuning and after prompt fine-tuning. We refer to the non-fine-tuned version as TWIN-GPT-origin. The accuracy of TWIN-GPT and TWIN-GPT-origin in severe outcome prediction is 0.821 and 0.537, respectively. Moreover, the attribute disclosure of TWIN-GPT when $x$ set to 20 is less than 0.3, but in contrast, TWIN-GPT-origin reaches 0.530, meaning that attackers can easily infer additional attributes of an individual based on the features of a subset of the data he knows. In terms of the indicator of Personalized Generation quality and Counterfactual Generation quality, most of the participants’ $r$ for TWIN-GPT are larger than 0.8, but the majority of participants’ $r$ for TWIN-GPT-origin are less than 0.2.
Metric TWIN-GPT TWIN-GPT-origin
Accuracy in Severe Outcome Prediction 0.821 0.537
Attribute Disclosure when $x=20$ 0.260 0.530
Personalized Generation ($r$ value) 0.812 0.121
Counterfactual Generation ($r$ value) 0.785 0.310
Table 1: Ablation study of TWIN-GPT and TWIN-GPT-origin